Session: Feb 2006 (session 1)
Wavelength: 5 cm


To give feedback on this experiment, click

Experiment Feedback

Torun (Kaz Borkowski at Wed Feb 22 12:50:50 2006) Failure : ANT
Due to severe snow and wind there was prolonged (some 6 hours) power down before this experiment start. Now with electricity on, our telescope still cannot be moved in one of the axles and cooling systems are in sorrowful condition. Works are going on but it seems it will take considerable time until we will be ready to continue observations.

Medicina (scalambra at Wed Feb 22 13:02:02 2006) Success :

Hartebeesthoek (Jonathan Quick at Wed Feb 22 13:38:48 2006) Success :
No known problems

Effelsberg (A. Polatidis at Wed Feb 22 15:09:32 2006) Minor failures :
Weather: Overcast

Onsala ( Michael Lindqvist at Wed Feb 22 15:27:48 2006) Success :
No known problems.

Westerbork (Tony Foley at Thu Feb 23 10:22:58 2006) Success :
Single dish was used. Please change the coordinates in the vexfile accordingly. No known problems

Jodrell_Bank (Alastair Gunn at Fri Mar 3 17:06:40 2006) Success :
The MkII telescope was used (Jb2). No known problems.

Cambridge (Alastair Gunn at Fri Mar 3 17:07:05 2006) Success :
No known problems.

Noto (P. Cassaro at Tue Mar 7 15:39:07 2006) Success :

EVN experiment feedback facility: entire session

Entire Session

Session: Feb 2006 (session 1)
Wavelengths: 6 3.6 5 18 cm


To give feedback on this entire session, click

Entire Session Feedback

Torun (Kaz Borkowski at Mon Feb 20 11:25:29 2006) Minor failures : PCAL
Tr phase cal system failed to produce signals during all the 6 cm experiments for yet unknown reason.

Westerbork (Tony Foley at Fri Feb 24 15:48:24 2006) Minor failures : ANT POIN
The feedback from the antenna control for SINGLE-DISH observations is unreliable for 20seconds after a source change. We will try to fix this. (Affected only the 5cm experiments -Z.P.)